a wheel makes 20 revolutions per second is stopped in 20s. its angular retardation is

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
as we know
 
w(t) = w(0) + alpha * t
 
hence
 
angular retardation alpha = 20/20 = 1 
 
hope it helps
 
Regards
Arun